OpenStudy (anonymous):

given a temperature 0f 300 kelvin, what is the approximate temperature in degrees Celsius?

OpenStudy (anonymous):

See K = 273 + C So, here K = 300 Therefore, 300 - 273 = C In Celcius, the temperature is : 27..

OpenStudy (anonymous):

your answer would be 27

OpenStudy (anonymous):

Valid answers (sig fig = # of significant figures): 1 sig fig = 30°C 2 sig fig = 27°C 4 sig fig (exact) = 26.85°C

OpenStudy (anonymous):

Thank You!